i do nto know how to answer sphere let alone find the radus is there any easy whay to find this
130.
Radius of a sphere
The radius of sphere is given by the formula r=(0.75v/pie)^1/3
Where v is it’s the volume. Find radius of a spherical tank that has a volume of 32pie/3 cubic meters

Answer by stanbon(75887)

r=(0.75v/pie)^1/3
Where v is it’s the volume. Find radius of a spherical tank that has a volume of 32pie/3 cubic meters
---------------------
r=(0.75/pi)^(1/3)
r=(0.75)(32/3)^(1/3)
r=(0.75)(2)(4/3)^(1/3)
r=(3)(4/3)^(1/3)
r=3.3019... meters
Cheers,
Stan H.